A 200.0 ml of a gas is collected at 20.0 0C and 733.5mm of Hg. The mass of the sample is 0.934g. What is its Molar mass?
1
Expert's answer
2020-03-30T07:57:19-0400

Clapeyron-Mendeleev equation:

"PV=\frac{m}{M}RT",

where "P" - pressure, atm; "V" - volume, l; "m" - mass of the gas, g; "M" - molar mass of the gas, g/mol; "R" - gas constant, "R=0.082 \; atm/(l^3 \cdot K)"; "T" - temperature, K.

"P[atm]=1.3158 \cdot 10^{−3} \cdot P [mm \; of \; Hg] = 1.3158 \cdot 10^{−3} \cdot 733.5 \approx 0.9651 \; atm"

"T[K]=273.15+T[\degree C]=273.15+20.0=293.15 \; K"

"M=\frac{mRT}{PV}"

"M=\frac{0.934 \cdot 0.082 \cdot 293.15}{0.9651 \cdot 200 \cdot 10^{-3}} \approx 116.32 \; g/mol"


Answer: 116.32 g/mol.